# Asynchronous Perception Machine for Efficient Test Time Training

### v3

@@ 1 Introduction
In these past centuries, computing-machines have become a lot faster [98]. This made it possible to train higher-parameterized neural nets and led to interesting emergent abilities [86]. As was predicted by Turing himself, and as were his suspicions of Lady Lovelace’s arguments against learning machines[96]: neural-nets can now finally learn without human-feedback [7], paint pictures [82] and even compose a sonnet [98]. Even with such impressive-progress, a key question still remains: how can these nets recognize images whose distribution is far different from the ones which were used during training? For e.g., consider a self-driving car trying to stop when it encounters a pedestrian crossing a road. Such practical scenarios require ‘instantaneous-decisions’ [^1] A flick of the eye, the fall of her hair, the passing of a second, or capturing the world at 120fps: instantaneous can be a subjective experience. Seconds can sometimes pass slowly for other reasons also. for ensuring human-safety in autonomous-systems [70].
@@
Figure 1: (i) Asynchronous Perception Machine (APM): An image $I$ passes through a column module and routes to a trigger column $T_{i}$. $T_{i}$ then unfolds and generates $h\times w$ location-specific queries. These queries are i.i.d and can be parallelized across cores of a gpu [53]. Each query $T_{i}$ is passed through a shared MLP and yields the vector $f^{\prime}_{xy}$ and $f^{\prime}_{rgb}$. MLP is queried iteratively until whole grid $f^{{}^{\prime}}$ comes into existence. Classification then involves comparing the averaged representation $f$ with class-specific textual representations in the contrastive space . (ii) Folded State: The parameters which the net learns are parameters of $T$ and MLP. (iii) Unfolded State: $T$ expands to yield $h\times w$ queries ‘on-the-fly’. Learning involves oscillating this net between folded and unfolded states. This net can be trained via backpropogation [84, 33].
@@ 2 Background
We draw from insights previously philosophically mentioned in GLOM [34]. Consider how machine perception has been done classically: an input $x\in\mathds{R}^{c\times h\times w}$ is transformed by a non-linear function $f$ to a perceptual feature grid $y\in\mathds{R}^{c^{\prime}\times h\times w}$, $c$ denotes image channels, $h,w$ are input spatial dimensions.
Islands of agreement [34] Rather than viewing this matrix $y$ as a cuboidal grid, one can now see this as column vectors $v_{c^{\prime}}$ at each location. There are $h\times w$ such columns in $y$. Therefore, there exists a one-to-one mapping, between each input patch i.e. $(x,y)$ with the vector $v_{c^{\prime}}$ at that location. A neural net $f$ can then thus learn to fit $f(I,x,y)\rightarrow v_{c^{\prime}}$. These vectors $v_{c^{\prime}}$ have been termed islands of agreement in GLOM [34]. They were recently demonstrated on realistic datasets by [70], showing some promise that trained models can serve as free supervision-sources for vectors $v_{c^{\prime}}$.
@@ 3.2 Global Column Module: Defining compressed representation T
We define a column $T$ as a vector of dimensions $1\times 1\times d$. Our aim is to map image $I$ in this $T$, so that $T$ can summarize its entire identity. Given an image of dimensions $c\times h\times w$, we run a 2D convolution on it. Number of filters are set as accordingly. The resultant $1\times h^{\prime}\times w^{\prime}$ feature map is then flattened into a single column vector T of dimensions $d=h^{\prime}\times w^{\prime}$ We shall refer to this column T as "triggering hyper-column"(seed). The only learnable parameters in this column are parameters of a convolution filter. [^2] The word triggering has been chosen because this column shall be used to trigger the queries to the neural field shared across different locations [34].[98, 97].
@@ 3.4 Computational Principle: Location-aware columns and their collapse
The unfolding-process shall now be concerned with generating location-aware column $T_{ij}$ from $T$. We generate $h*w$ 2-D non-parametric positional encoding similar to the ones being used in transformers [100] and neural fields [65]. The trigger column $T_{ij}$ is then given by $T_{ij}=(T|p_{ij})$, where |||| denotes the stacking operator and $(1,1)\leq(i,j)\leq(H,W)$. T can be said to encode identity of an image.
@@ 4 Experimenting with APM
Results and Analysis: We study APM’s performance on test-time-training on several datasets. APM processes each test sample individually: i.e. the weights are drawn from a normal distribution after processing every sample to prevent information leakage. For zero-shot classification of a test-sample, APM leverages the $80$ textual-prompt templates similar to the ones used in CLIP. In Tab 1, APM scales up to zero-shot classification task to datasets with $1000$ classes. Using CLIP-ViT B/16 as a teacher, we surpass TPT [90] with an avg score of $62.6$ and avg ood-score of $61.2$. Next, we benchmark OpenCLIP-VITH/14 against all these splits. Using the same model as our teacher, we get an absolute improvement of $3\%\uparrow$ $[84.6\%]$ on ImageNet val set, $5.1\%\uparrow$ $[84.2\%]$ on ImageNet-A, $3.2\%\uparrow$ $[83.9\%]$ on ImageNet-V2, $2\%\uparrow$ $[94.9\%]$ on ImageNet-R, $4.3\%\uparrow$ $[77.1\%]$ on ImageNet-Sketch respectively. A similar trend is observed with a VIT-L/14 backbone. This might lead to the conclusion that a stronger teacher seems to benefit APM. In Table 2, we show results on Imagenet-C. Using a CLIP VIT-L/14 teacher, APM gets highest accuracy on $11/15$ noises with the highest average score of $50.3$. Note that TTT-MAE also uses a VIT-L encoder, and is $pre\-trained$. In contrast, $APM$ doesn’t need any dataset specific-pretraining. Finally, in Tab LABEL:tab:fine-grained, APM improves upon $4/9$ datasets, comes close on remaining $5$ and gets the highest average accuracy score of $65.5$.
@@
Table 4: APM’s computational analysis: TTT for 20 iterations on APM. Baseline is CLIP VIT-B/16 which is used as a teacher in [90]. $M_{meas}(GB)$, $GFlops_{meas}(GB)$ are tmeasured stats. $M_{i}(GB)$, $GFlops_{i}(GB)$ are idealistic estimates. (t): $t^{th}$ ttt iteration, (s): student, (u): teacher, (s+u): portion of memory/flops consumed by student/teacher respectively. Note that APM is a $25M$ net.
$t$ Params(M)$\downarrow$ $M_{meas}(GB)$downarrow $M_{i}(GB)$downarrow $GFlops_{meas}$downarrow $GFlops_{i}$downarrow Swin[57] 1-20 87 1.5 1.4 353 308 TPT[90] 1-20 151.3 3.1 2.7 529 476 CLIP VIT-B/16 1-20 149.2 2.3 1.8 462 410 CLIP VIT-B/16(u)[90] 1 149.2 1.8 1.8 20.5 20.5 APM(s) 1 174.2(s+u) 2.7 (s+u) 1.8(u) + 0.6(s) 20.5(u) 20.5(u) APM(s) 2 174.2(s+u) 2.7 (s+u) 1.8(u) + 0.6(s) 10(s) 10(s) Peak Use 1-20 174.2(s+u) 2.7 (s+u) 1.8(u) + 0.6(s) 241.7 (s+u) 210.5 (s+u)
@@ 5 APM Training (Qualitative Analysis)
APM can also scale up and do learning on a batch of samples (for e.g., COCO images[55]) distilled from a teacher[36]. This requires introducing several new mechanisms. Note that this section is meant to qualitatively demonstrate how scaling up APM can improve the net’s interpretability, and help seed future research. Quantitative experiments beyond test-time-training remain out of the scope of this paper. APM’s training follows a standard setup in self-supervised-learning[7]. We have provided the full algorithm for SSL-training/test-time-training in the AppendixC. During inference, APM takes any 2D image $x_{k}$ and predicts its RGB reconstruction $RGB{k}$/higher dimensional features $f_{k}$. The net then begins to demonstrate several interesting properties, which we will discuss next.
@@
Given a sample $x_{k}$, APM can reconstruct its RGB. In Fig4, we achieve this by estimating $f_{rgb,ij}=(T_{ij}|f_{ij})$. This skip-connection from trigger column $T_{ij}$ to output feature has a subtle reason: consider a white dog and brown dog. The predicted object-level feature for both will be almost identical[34, 2]. However, $T_{ij}$ is different for both since it contains lower patch-level features[34]. Therefore, this helps us break symmetry. Without this skip-connection[25], the net fails to predict RGB. The network is trained to reconstruct RGB for a batch of images, $L_{rgb}=\sum_{i=1}^{N}\sum_{j=1}^{h*w}L_{2}(p^{\prime}_{j},p_{j})$, where $p_{j}$/$p_{j}^{{}^{\prime}}$ are ground truth/predicted RGB-logits respectively.
@@ 6 Ablations on APM
(b) Ablations of our APM on C-10.$I_{xy}$ means that local patch was injected into the column $T_{c}$.
@@ 8 Conclusion
We propose APM, inspired from the insights presented in GLOM[34, 43]. APM promises to be an efficient architecture for test-time-training and asynchronous patch-processing[97]. APM shows robustness to extreme distribution shifts[28]. APM demonstrates that MLP’s can be made to semantically cluster a given image. We hope that APM will help inspire further research on simpler weight-sharing, lower-memory, higher-bandwidth baby-friendly [^6] Thku thku so much for tracks for babies like us[23]. efficient-nets [33]. We love GLOM shoo much and everyone who keeps helping it. [34].
little_dogzilla [Uncaptioned image]
@@ Appendix B Future Work
              /* Route sample $x_{k}$ in column $T$[34] */
@@
              /* Route sample $x_{k}$ in column $T$[34] */
@@
              /* Collapse all location-specific queries $T_{ij}$ into $T$[34] */
@@ E.1 Architecture and Hyperparameters
Architecture: We inflate APM’s architecture in Table6. For the TTT experiments, APM consists of only a single convolutional layer, and 5 MLP layers. Additionally, APM consists of a feature projection head containing a single linear layer, and an optional RGB head. It maybe noted that the number of kernels in the convolutional filter is only 1. This creates a subtle issue: the RGB reconstruction in Fig4 is black and white. This is because a single kernel loses RGB channel information. Put simply, assume a tuple of 3 numbers representing RGB values, <1,2,3><1,2,3>< 1 , 2 , 3 > ,<4,1,1><4,1,1>< 4 , 1 , 1 >. For a convolution operation with a single kernel assuming unit weights, the answer is $6$. If this $6$ gets injected in the net, it is equally certain that the input was <1,2,3><1,2,3>< 1 , 2 , 3 > or a <4,1,1><4,1,1>< 4 , 1 , 1 > making reconstruction from the RGB head of APM difficult. We found that this symmetry-breaking issue could be noticeably fixed with $n_{kernels}\geq n_{channels}$, where $n_{channels}$ is the number of channels $c$ in the input. Historically, various other rotational/translational/mirror-symmetries have played an important role in designing boltzmann machines[88]. The architecture in Table 6 is then meant to showcase APM’s potential to an extrema: how much can it do even with a single convolutional filter?
@@
Decoder Linear $(d_{p}+d)$ $4096$
Linear $4096$ $4096$
Linear $4096$ $4096$
Linear $4096$ $2048$
Linear $2048$ $1024$
Feature Projection Head Linear $1024$ $d_{c}$
RGB-Head(optional) Linear $(d_{p}+d+1024)$ $4096$
Linear $1024*3$ $256$
Linear $256$ $256$
Linear $256$ $3$
@@ Appendix G Additional Results
Using it as a teacher, we get the lowest average error rate of $14.8\%$, thereby even improving upon the performance of CLIP VIT-L/14. Note that our APM is reinitialized with random weights after every-test sample in contrast to methods like TTT-Online which $retain$ the weights after every ttt-iteration. Inspite of that, we get a lower error rate i.e. $14.8\%$ than TTT-Online $19.1\%$. Another benefit which APM gains is that it can directly use the textual-encoder of the Clip VIT-L/14 teacher to classify on cifar-10 test set: this allows us to bypass the requirement of training a separate dataset-specific linear probe on top of our net.
@@ Appendix H Some helpful analogies
A computational analogy[34]: We now start treating an image $I$ as a digital organism. It starts from some compressed representation $T$. $T$ unfolds to yield the image $I$. $I$ then folds back to yield the compressed representation $T$. Learning proceeds by oscillating between these unfolded and folded phases. At every step, the net is trying to reconstruct image $I$ from $T$. $T$ is then expected to be a dense vector-space.
@@ Appendix I A new representation: Hinton’s Islands of agreement[70, 34]
We present the islands of agreement which can be observed now in APM in Fig 7,10,11. Notice how the part-wholes in the images are clearly observable in different colours. Finally, we scaled up our APM and trained it on COCO train set. Fig 13,12,8,14 visualizes the islands of agreement on COCO val set. It can clearly be seen that the net develops a semantic clustering and shows the potential to do a dense task. APM offers a unique advantage over parallel perception (DINOv2): features at a particular location can be queried serially. One can ‘selectively’ query the locations which correspond to a particular object[109], instead of bringing the whole feature grid in existence and then choosing the relevant objects[6]. Such an inductive bias of choosing which of the input rays/columns in a neural field corresponds to which object has already been explored by Yu et al[109].
@@
Quod errat demonstrandum. Q.E.D[15]. May these humble neural-nets[66] now fold themselves[31] and stay safe[4]. For all of us. The last section of forward-forward paper talks about Mortal computation and it now appears that we might reach that point[33]. AI safety however still needs a lot more work[27].
Till the next time the space-curves[13],
little__\__dogzilla aboard ‘Starship Connectionist Enterprise’.
End of transmission.
Figure 9: A cute-little-godzilla trying its best to navigate wormholes [104]. Hiyaaa!!

### v4

@@ 1 Introduction
In these past centuries, computing-machines have become a lot faster [92]. This made it possible to train higher-parameterized neural nets and led to interesting emergent abilities [80]. As was predicted by Turing himself, and as were his suspicions of Lady Lovelace’s arguments against learning machines[90]: neural-nets can now finally learn without human-feedback [6], paint pictures [76] and even compose a sonnet [92]. Even with such impressive-progress, a key question still remains: how can these nets recognize images whose distribution is far different from the ones which were used during training? For e.g., consider a self-driving car trying to stop when it encounters a pedestrian crossing a road. Such practical scenarios require ‘instantaneous-decisions’ for ensuring human-safety in autonomous-systems [64].
@@
Figure 1: (i) Asynchronous Perception Machine (APM): An image II passes through a column module and routes to a trigger column $T_{i}$. $T_{i}$ then unfolds and generates $h\times w$ location-specific queries. These queries are i.i.d and can be parallelized across cores of a gpu [48]. Each query $T_{i}$ is passed through a shared MLP and yields the vector $f^{\prime}_{xy}$ and $f^{\prime}_{rgb}$. MLP is queried iteratively until whole grid $f^{{}^{\prime}}$ comes into existence. Classification then involves comparing the averaged representation $f$ with class-specific textual representations in the contrastive space . (ii) Folded State: The parameters which the net learns are parameters of $T$ and MLP. (iii) Unfolded State: $T$ expands to yield $h\times w$ queries ‘on-the-fly’. Learning involves oscillating this net between folded and unfolded states. This net can be trained via backpropogation [78, 28].
@@ 2 Background
We draw from insights previously philosophically mentioned in GLOM [29]. Consider how machine perception has been done classically: an input $x\in\mathds{R}^{c\times h\times w}$ is transformed by a non-linear function $f$ to a perceptual feature-cuboid $y\in\mathds{R}^{c^{\prime}\times l\times h\times w}$, $c$ denotes image channels, $l$ are the intermediate layers, $h,w$ are input spatial dimensions. Note that $y$ is estimated after feed-forward through $l$ layers.
@@ 3.2 Global Column Module: Defining compressed representation T
We define a column $T$ as a vector of dimensions $1\times 1\times d$. Our aim is to map image $I$ in this $T$, so that $T$ can summarize its entire identity. Given an image of dimensions $c\times h\times w$, we run a 2D convolution on it. Number of filters are set as accordingly. The resultant $1\times h^{\prime}\times w^{\prime}$ feature map is then flattened into a single column vector T of dimensions $d=h^{\prime}\times w^{\prime}$ . We shall refer to this column T as "triggering hyper-column"(seed). The only learnable parameters in this column are parameters of a convolution filter. [^1] The word triggering has been chosen because this column shall be used to trigger the queries to the neural field shared across different locations [29].[92, 91].
@@ 3.4 Computational Principle: Location-aware columns and their collapse
The unfolding-process shall now be concerned with generating location-aware column $T_{ij}$ from $T$. We generate hwh*w 2-D non-parametric positional encoding similar to the ones being used in transformers [94] and neural fields [60]. The trigger column $T_{ij}$ is then given by $T_{ij}=(T|p_{ij})$, where $|$ denotes the stacking operator and $(1,1)\leq(i,j)\leq(H,W)$. T can be said to encode identity of an image.
@@ 4 Experimenting with APM
Results and Analysis: We study APM’s performance on test-time-training on several datasets. APM processes each test sample individually: i.e. the weights are drawn from a normal distribution after processing every sample to prevent information leakage. For zero-shot classification of a test-sample, APM leverages the $80$ textual-prompt templates similar to the ones used in CLIP. In Tab 1, APM scales up to zero-shot classification task to datasets with $1000$ classes. Using CLIP-ViT B/16 as a teacher, we surpass TPT [84] with an avg score of $62.6$ and avg ood-score of $61.2$. Next, we benchmark OpenCLIP-VITH/14 against all these splits. Using the same model as our teacher, we get an absolute improvement of $3\%\uparrow$ $84.6\%]$ on ImageNet val set, $5.1\%\uparrow$ $84.2\%]$ on ImageNet-A, $3.2\%\uparrow$ $83.9\%]$ on ImageNet-V2, $2\%\uparrow$ $94.9\%]$ on ImageNet-R, $4.3\%\uparrow$ $77.1\%]$ on ImageNet-Sketch respectively. A similar trend is observed with a VIT-L/14 backbone. This might lead to the conclusion that a stronger teacher seems to benefit APM. In Table 2, we show results on Imagenet-C. Using a CLIP VIT-L/14 teacher, APM gets highest accuracy on $11/15$ noises with the highest average score of $50.3$. Note that TTT-MAE also uses a VIT-L encoder, and is pretrainedpre\-trained. In contrast, $APM$ doesn’t need any dataset specific-pretraining. Finally, in Tab LABEL:tab:fine-grained, APM improves upon $4/9$ datasets, comes close on remaining $5$ and gets the highest average accuracy score of $65.5$.
@@
Table 4: APM’s computational analysis: TTT for 20 iterations on APM. Baseline is CLIP VIT-B/16 which is used as a teacher in [84]. $M_{meas}(GB)$, $GFlops_{meas}(GB)$ are tmeasured stats. $M_{i}(GB)$, $GFlops_{i}(GB)$ are idealistic estimates. (t): $^{th} tt$t iteration, (s): student, (u): teacher, (s+u): portion of memory/flops consumed by student/teacher respectively. Note that APM is a $25M$ net.
[…]
@@ 5 APM Training (Qualitative Analysis)
APM can also scale up and do learning on a batch of samples (for e.g., COCO images[50]) distilled from a teacher[31]. This requires introducing several new mechanisms. Note that this section is meant to qualitatively demonstrate how scaling up APM can improve the net’s interpretability, and help seed future research. Quantitative experiments beyond test-time-training remain out of the scope of this paper. APM’s training follows a standard setup in self-supervised-learning[6]. We have provided the full algorithm for SSL-training/test-time-training in the AppendixC. During inference, APM takes any 2D image $x_{k}$ and predicts its RGB reconstruction RGBkRGB{k}/higher dimensional features $f_{k}$. The net then begins to demonstrate several interesting properties, which we will discuss next.
@@
Given a sample $x_{k}$, APM can reconstruct its RGB. In Fig4, we achieve this by estimating $f_{rgb,ij}=(T_{ij}|f_{ij})$. This skip-connection from trigger column $T_{ij}$ to output feature has a subtle reason: consider a white dog and brown dog. The predicted object-level feature for both will be almost identical[29, 2]. However, $T_{ij}$ is different for both since it contains lower patch-level features[29]. Therefore, this helps us break symmetry. Without this skip-connection[22], the net fails to predict RGB. The network is trained to reconstruct RGB for a batch of images, $L_{rgb}=\sum_{i=1}^{N}\sum_{j=1}^{h*w}L_{2}(p^{\prime}_{j},p_{j})$, where $p_{j}$$p_{j}^{{}^{\prime}}$ are ground truth/predicted RGB-logits respectively.
@@ 6 Ablations on APM
[…]
@@ 8 Conclusion
We propose APM, inspired from the insights presented in GLOM[29, 38]. APM promises to be an efficient architecture for test-time-training and asynchronous patch-processing[91]. APM shows robustness to extreme distribution shifts[24]. APM demonstrates that MLP’s can be made to semantically cluster a given image. We hope that APM will help inspire further research on simpler weight-sharing, lower-memory, higher-bandwidth efficient-nets [28].
@@ Appendix B Future Work
    /* Route sample $x_{k}$ in column TT[29] */
@@
    /* Route sample $x_{k}$ in column TT[29] */
@@
    /* Collapse all location-specific queries $T_{ij}$ into TT[29] */
@@ E.1 Architecture and Hyperparameters
Architecture: We inflate APM’s architecture in Table6. For the TTT experiments, APM consists of only a single convolutional layer, and 5 MLP layers. Additionally, APM consists of a feature projection head containing a single linear layer, and an optional RGB head. It maybe noted that the number of kernels in the convolutional filter is only 1. This creates a subtle issue: the RGB reconstruction in Fig4 is black and white. This is because a single kernel loses RGB channel information. Put simply, assume a tuple of 3 numbers representing RGB values, <1,2,3><1,2,3> ,<4,1,1><4,1,1>. For a convolution operation with a single kernel assuming unit weights, the answer is $6$. If this $6$ gets injected in the net, it is equally certain that the input was <1,2,3><1,2,3> or a <4,1,1><4,1,1> making reconstruction from the RGB head of APM difficult. We found that this symmetry-breaking issue could be noticeably fixed with $n_{kernels}\geq n_{channels}$, where $n_{channels}$ is the number of channels $c$ in the input. Historically, various other rotational/translational/mirror-symmetries have played an important role in designing boltzmann machines[82]. The architecture in Table 6 is then meant to showcase APM’s potential to an extrema: how much can it do even with a single convolutional filter?
@@
Decoder Linear $(d_{p}+d)$ * $4096$ - - -
Linear $4096$ * $4096$ - - -
Linear $4096$ * $4096$ - - -
Linear $4096$ * $2048$ - - -
Linear $2048$ * $1024$ - - -
Feature Projection Head Linear $1024$ * $d_{c}$ - - -
RGB-Head(optional) Linear $(d_{p}+d+1024)$ * $4096$ - - -
Linear 102431024*3 * $256$ - - -
Linear $256$ * $256$ - - -
Linear $256$ * 33 - - -
@@ Appendix G Additional Results
Using it as a teacher, we get the lowest average error rate of $14.8\%$, thereby even improving upon the performance of CLIP VIT-L/14. Note that our APM is reinitialized with random weights after every-test sample in contrast to methods like TTT-Online which retainretain the weights after every ttt-iteration. Inspite of that, we get a lower error rate i.e. $14.8\%$ than TTT-Online $19.1\%$. Another benefit which APM gains is that it can directly use the textual-encoder of the Clip VIT-L/14 teacher to classify on cifar-10 test set: this allows us to bypass the requirement of training a separate dataset-specific linear probe on top of our net.
@@ Appendix H Some helpful analogies
A computational analogy[29]: We now start treating an image $I$ as a digital organism. It starts from some compressed representation $T$. $T$ unfolds to yield the image $I$. II then folds back to yield the compressed representation $T$. Learning proceeds by oscillating between these unfolded and folded phases. At every step, the net is trying to reconstruct image $I$ from $T$. $T$ is then expected to be a dense vector-space.
@@ Appendix I A new representation: Hinton’s Islands of agreement[64, 29]
We present the islands of agreement which can be observed now in APM in Fig 7,9,10. Notice how the part-wholes in the images are clearly observable in different colours. Finally, we scaled up our APM and trained it on COCO train set. Fig 12,11,8,13 visualizes the islands of agreement on COCO val set. It can clearly be seen that the net develops a semantic clustering and shows the potential to do a dense task. APM offers a unique advantage over parallel perception (DINOv2): features at a particular location can be queried serially. One can ‘selectively’ query the locations which correspond to a particular object[102], instead of bringing the whole feature grid in existence and then choosing the relevant objects[5]. Such an inductive bias of choosing which of the input rays/columns in a neural field corresponds to which object has already been explored by Yu et al[102].